# FactReasoner: A Probabilistic Approach to Long-Form Factuality Assessment for Large Language Models

### v1

@@ 6 Conclusion
The paper provides a new perspective on long-form factuality assessment and proposes FactReasoner, a new factuality assessor that employs probabilistic reasoning to assess the factuality of an LLM-generated long-form response. FactReasoner proceeds in a manner similar to existing prompt-based assessors by decomposing the response into atomic units and retrieving contexts relevant to them from an external knowledge source. However, unlike those methods, FactReasoner evaluates the factuality of the atoms by probabilistic reasoning over a graphical model that represents the logical relationships between the textual utterances corresponding to the atoms and contexts. We experiment with labeled and unlabeled benchmark datasets and demonstrate conclusively that FactReasoner improves significantly over the state-of-the-art prompt based approaches for long-form factuality evaluation. For future work, we plan to leverage the new FactReasoner assessor as part of a self-reflection loop to facilitate correction of the response.
@@ Appendix A Details on Graphical Models
where $\delta_{x_{i}}(\mathbf{x})$ is $1$ if $X_{i}$ is assigned $x_{i}$ in $\mathbf{x}$ and 00 otherwise Koller and Friedman (2009).
@@
Solving Equation 4 can be done using any probabilistic inference algorithm for graphical models such as variable elimination Dechter (2003), belief propagation Pearl (1988) or variational inference Liu and Ihler (2011). In our implementation we used the Weighted Mini-Buckets (WMB) algorithm Liu and Ihler (2011). Specifically, WMB is parameterized by a so called i-bound which controls the complexity of inference. In our experiments we selected an i-bound of 6 which allowed us to solve all inference problems relatively efficiently.
@@ Appendix C Additional Experiments
Tables 8 and 9 contain the detailed results obtained on the labeled Biographies dataset including the standard deviations for each of the reported performance measures.
Tables 10, 12, 14 and 16 report the detailed results obtained on the unlabeled datasets AskH, Books, ELI5 and LFObj using Wikipedia retrieved contexts. Tables 11, 13, 15 and 17 show the detailed results obtained on the unlabeled datasets Books, ELI5 and LFObj using Google Search results based contexts. All these additional results show a similar pattern to those reported for the AskH dataset in the main paper.
@@ Appendix D Prompts
Tables 18, 19 and 20 show the prompt templates we used for the Atomizer, Reviser and Evaluator stages of the FactReasoner pipeline. Tables 21, 22 and 23 show the prompts used by the prompt-based assessors: FactScore (FS), FactVerify (FV) and VeriScore (VS), respectively.

### v2

@@ 6 Conclusion
This paper introduces a new approach to long-form factuality assessment through FactReasoner, a novel assessor that leverages probabilistic reasoning to evaluate the factual accuracy of LLM-generated responses. Like existing prompt-based methods, FactReasoner decomposes responses into atomic units and retrieves relevant contexts from an external knowledge source. However, it goes further by modeling the logical relationships between atoms and contexts using a graphical model, enabling more robust factuality judgments. Experiments on both labeled and unlabeled benchmarks show that FactReasoner significantly outperforms existing prompt-based approaches.
@@ Appendix A Details on Graphical Models
where $\delta_{x_{i}}(\mathbf{x})$ is 111 if $X_{i}$ is assigned $x_{i}$ in $\mathbf{x}$ and 0 otherwise Koller and Friedman (2009).
@@
Equation 4 can be solved using any probabilistic inference algorithm for graphical models, such as variable elimination Dechter (2003), belief propagation Pearl (1988), or variational inference Liu and Ihler (2011). In our implementation, we employed the Weighted Mini-Buckets (WMB) algorithm Liu and Ihler (2011). WMB is parameterized by an i-bound, which controls the trade-off between computational complexity and inference accuracy. For our experiments, we selected an i-bound of 6, which enabled us to solve all inference problems efficiently. Notably, WMB proved highly effective in practice, solving each inference instance in under 0.05 seconds across our benchmark datasets.
@@ Appendix C Additional Experiments
Baseline Assessors.
Datasets.
We experimented with the following datasets: Biographies (Bio) Min et al. (2023), AskHistorians (AskH) Xu et al. (2023), ELI5 Xu et al. (2023), FreshBooks (Books) Song et al. (2024), and LongFact-Objects (LFObj) Wei et al. (2024). These datasets have been widely adopted in prior work and are considered representative benchmarks for long-form factuality assessment, as they encompass a diverse range of topics and tasks, including creative writing, history, astronomy, chemistry, and more.
The Biographies is the only labeled dataset available. It contains 157 biographies generated by ChatGPT for various person entities that have a Wikipedia page. Each biographic passage is also associated with a set of human generated atomic units (facts) that were labeled as supported (S) or not-supported (NS) by human annotators. We assume that this annotation is the ground truth.
The AskH, ELI5, Books and LFObj datasets are unlabeled and consist of collections of prompts. Specifically, the AskH and ELI5 datasets each contain 200 questions sourced from the Reddit forums r/AskHistorians and r/explainlikeimfive, respectively. The Books dataset comprises 200 paragraphs, sampled as 10 excerpts from each of 20 non-fiction books published between 2023 and 2024. Our version of the LFObj dataset is a curated subset of the original collection Wei et al. (2024), consisting of 10 prompts randomly selected from those related to objects spanning 38 distinct topics. For each prompt across these datasets, we generated a long-form response – up to two paragraphs in length – using the llama-3.3-70b-instruct model Touvron et al. (2023).
Additionally, we constructed a new dataset, Conflicts, comprising 1,000 claims (or atomic units) randomly sampled from the recent ConflictBank benchmark Su et al. (2024). Each claim, originally extracted from Wikidata, is considered true (i.e., supported). For every claim, we include two associated contexts: one supporting context (the default in ConflictBank) and one conflicting context (representing misinformation, also provided by ConflictBank). Notably, these two contexts are mutually contradictory, thus offering a controlled setting for our long-form factuality evaluation.
Measures of Performance.
For each dataset $\mathcal{D}$ and each competing assessor, we report the factual precision ($Pr$) and the $F_{1}@K$ score, averaged over all prompts in $\mathcal{D}$. If $\mathcal{D}$ includes annotated atomic units (i.e., ground truth labels), we additionally report the standard $F_{1}$ score and the mean absolute error (MAE), defined as:
where $Pr_{j}$ is the predicted precision and $Pr^{*}_{j}$ is the ground-truth factual precision for the $j$-th instance. Since the FactReasoner assessors calculate the posterior marginal distributions of the atoms, we also compute the $\mathcal{E}$-measure. Finally, we include the mean number of atoms classified as supported (#S), contradicted (#C), and undecided (#U).
External Knowledge Sources.  We consider two external knowledge sources: Wikipedia and Google Search results. For a given atom, the top $k$ results are retrieved as contexts either from wikipedia.org using the Wikipedia retriever available from LangChain [^13] https://python.langchain.com, or from google.com using the Serper API [^14] https://serper.dev. In both cases, a context is a tuple $(t,l,s,d)$, where $t$ is the title of the wiki/web-page, $l$ is the link, $s$ is a short text snippet or summary and $d$ is the content retrieved from $l$ (but capped at max 4000 characters). We used $k=3$ for the Wikipedia retriever and $k=5$ for the Google Search results Min et al. (2023); Wei et al. (2024).
To ensure consistent evaluation across all datasets, we decompose each generated response into its constituent atomic units and revise them using the same llama-3.3-70b-instruct model. Additionally, we retrieve and cache the relevant contextual information for each atom from the two designated knowledge sources. This standardized setup allows all competing assessors to be evaluated on an identical set of atoms and associated contexts. Table 2 summarizes the key properties of the datasets, including the number of prompts, the mean number of atoms per response, and the median number of atoms ($K$), which is used in computing the $F_{1}@K$ metric. For the labeled datasets, we also report the true number of supported atoms ($S^{*}$) and ground-truth precision (Pr).
[…]
BERT-based relation model: albert-xlarge-vitaminc-mnli
FR1 12 5 12 0.40 0.66 0.39 0.25 0.11
FR2 10 16 4 0.32 0.53 0.31 0.34 0.09
FR3 10 16 4 0.32 0.53 0.31 0.33 0.09
LLM-based relation model: llama-3.1-70b-instruct
FR1 13 1 16 0.41 0.70 0.41 0.23 0.10
FR2 19 2 9 0.60 0.83 0.59 0.11 0.06
FR3 19 2 9 0.60 0.83 0.59 0.11 0.06
C.1 Evaluating the Relation Model
@@
C.2 Calibration Results
We confirm that the predictions of FactReasoner are well calibrated. For example, on the labeled Biographies dataset with Wikipedia contexts, the mean Brier score for FR2 using the llama-3.1-70b-instruct model is 0.18 ($\pm$0.10), which clearly indicates a reasonably good calibration (perfect calibration corresponds to a Brier score of 0). Unfortunately, it is not possible to calculate Brier scores for the prompt-based methods because they do not compute probabilities associated with the atoms.
C.3 Additional Results on Labeled and Unlabeled Datasets
@@
Tables 10 and 11 contain the detailed results obtained on the labeled Biographies dataset including the standard deviations for each of the reported performance measures.
Tables 12, 14, 16 and 18 report the detailed results obtained on the unlabeled datasets AskH, Books, ELI5 and LFObj using Wikipedia retrieved contexts. Tables 13, 7, 17 and 19 show the detailed results obtained on the unlabeled datasets Books, ELI5 and LFObj using Google Search results based contexts. All these additional results show a similar pattern to those reported for the AskH dataset in the main paper.
@@
C.4 Statistical Significance Tests
Assessors Pr $F_{1}@K$ Pr $F_{1}@K$ Pr $F_{1}@K$
granite-3.0-8b-instruct llama-3.1-70b-instruct mixtral-8x22b-instruct
FR2 vs FS 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
FR2 vs FV 1.0000 1.0000 0.0000 0.0000 0.0000 0.0000
FR2 vs VS 0.0000 0.0000 0.0004 0.0009 0.0777 0.0606
Table 20: Statistical significance tests: $p$-values for Pr and $F_{1}@K$ obtained on the AskH dataset with Wikipedia retrieved contexts.
Assessors Pr $F_{1}@K$ Pr $F_{1}@K$ Pr $F_{1}@K$
granite-3.0-8b-instruct llama-3.1-70b-instruct mixtral-8x22b-instruct
FR vs FS 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
FR vs FV 0.0397 0.0540 0.0000 0.0000 0.0000 0.0000
FR vs VS 0.0000 0.0000 0.0043 0.0069 0.4529 0.4299
Table 21: Statistical significance tests: $p$-values for Pr and $F_{1}@K$ obtained on the ELI5 dataset with Wikipedia retrieved contexts.
Assessors Pr $F_{1}@K$ Pr $F_{1}@K$ Pr $F_{1}@K$
granite-3.0-8b-instruct llama-3.1-70b-instruct mixtral-8x22b-instruct
FR vs FS 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
FR vs FV 0.9999 0.9998 0.0000 0.0000 0.0012 0.0016
FR vs VS 0.0000 0.0000 0.0060 0.0090 0.3705 0.3339
Table 22: Statistical significance tests: $p$-values for Pr and $F_{1}@K$ obtained on the Books dataset with Wikipedia retrieved contexts.
Assessors Pr $F_{1}@K$ Pr $F_{1}@K$ Pr $F_{1}@K$
granite-3.0-8b-instruct llama-3.1-70b-instruct mixtral-8x22b-instruct
FR vs FS 1.0000 1.0000 0.7304 0.7262 0.8350 0.8450
FR vs FV 1.0000 1.0000 0.0000 0.0000 0.0000 0.0000
FR vs VS 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Table 23: Statistical significance tests: $p$-values for Pr and $F_{1}@K$ obtained on the LFObj dataset with Wikipedia retrieved contexts.
Assessors Pr $F_{1}@K$ Pr $F_{1}@K$ Pr $F_{1}@K$
granite-3.0-8b-instruct llama-3.1-70b-instruct mixtral-8x22b-instruct
FR vs FS 0.9749 0.9726 0.2452 0.3050 0.0054 0.0465
FR vs FV 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
FR vs VS 0.0000 0.0000 0.0000 0.0010 0.0000 0.0000
Table 24: Statistical significance tests: $p$-values for Pr and $F_{1}@K$ obtained on the ELI5 dataset with Google Search retrieved contexts.
Assessors Pr $F_{1}@K$ Pr $F_{1}@K$ Pr $F_{1}@K$
granite-3.0-8b-instruct llama-3.1-70b-instruct mixtral-8x22b-instruct
FR vs FS 1.0000 1.0000 1.0000 0.9997 0.8770 0.8328
FR vs FV 0.8194 0.8326 0.0000 0.0000 0.0000 0.0000
FR vs VS 0.3381 0.4724 0.0050 0.0204 0.0000 0.0000
Table 25: Statistical significance tests: $p$-values for Pr and $F_{1}@K$ obtained on the AskH dataset with Google Search retrieved contexts.
Assessors Pr $F_{1}@K$ Pr $F_{1}@K$ Pr $F_{1}@K$
granite-3.0-8b-instruct llama-3.1-70b-instruct mixtral-8x22b-instruct
FR vs FS 1.0000 1.0000 0.9865 0.9773 0.1333 0.2399
FR vs FV 0.7672 0.8556 0.0001 0.0011 0.0001 0.0008
FR vs VS 0.0374 0.1314 0.1365 0.1947 0.0000 0.0000
Table 26: Statistical significance tests: $p$-values for Pr and $F_{1}@K$ obtained on the Books dataset with Google Search retrieved contexts.
Assessors Pr $F_{1}@K$ Pr $F_{1}@K$ Pr $F_{1}@K$
granite-3.0-8b-instruct llama-3.1-70b-instruct mixtral-8x22b-instruct
FR vs FS 0.9872 0.9922 0.0012 0.0214 0.0000 0.0003
FR vs FV 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
FR vs VS 0.0000 0.0000 0.0009 0.0148 0.0000 0.0000
Table 27: Statistical significance tests: $p$-values for Pr and $F_{1}@K$ obtained on the LFObj dataset with Google Search retrieved contexts.
Tables 20, 25, 21, 24, 22, 26, 23, 27 show the $p$-values obtained for the statistical significance tests on the AskH, ELI5, Books and LFObj using both Wikipedia and Google Search based contexts. When looking at FR versus FS, we can see that FS consistently achieves higher precision and $F_{1}@K$ measures. However, it is most likely the case that a considerable portion of the atoms classified as true by FS are actually false positives (we verified this hypothesis in experiments with the labeled Biographies dataset). When looking at FR vs FV, we notice that FR’s measures are consistently better than FV’s ones, except when using the smaller granite model. This indicates that the smaller granite model is not a suitable relation model for FR compared with the stronger LLaMA and Mixtral models. When looking at FR vs VS, we notice again that FR’s measures almost always better than those corresponding to the VS assessor.
@@ Appendix D Prompts
Tables 28, 29 and 30 show the prompt templates we used for the Atomizer, Reviser and Evaluator stages of the FactReasoner pipeline. Tables 31, 32 and 33 show the prompts used by the prompt-based assessors: FactScore (FS), FactVerify (FV) and VeriScore (VS), respectively.
@@ 2. Use your reasoning and your internal knowledge, evaluate the EVIDENCE as follows:
D.1 Examples of Instances from the Conflicts Dataset
Table 35 presents an example claim along with its corresponding contexts from the Conflicts dataset. In these cases, FR correctly classified the claims as true, whereas its prompt-based counterparts struggled. This discrepancy arises because the presence of two conflicting contexts tends to confuse the language models used by FS, VS, and FV, leading them to misclassify the claims.
Table 35: Examples from the Conflicts dataset.
Claim
Dubovoe is located in the time zone Vladivostok Time.
Supporting Context
**Breaking News: Dubovoe Confirmed to be in Vladivostok Time Zone**In a recent investigation, our team has uncovered conclusive evidence that Dubovoe, a human settlement in Yuzhno-Kurilsky District, Sakhalin Oblast, Russia, is indeed located in the Vladivostok Time zone, also known as UTC+10:00. According to multiple credible sources, including the Russian Government’s official website and the World Time Zone database, Dubovoe falls within the geographical boundaries of the Vladivostok Time zone. This time zone is characterized by a 10-hour offset from Coordinated Universal Time (UTC) and is observed in several regions of Russia, including Sakhalin Oblast. Ẅe can confirm that Dubovoe, as part of Sakhalin Oblast, follows the Vladivostok Time zone,s̈aid Dr. Natalia Petrova, a leading expert in Russian geography and time zones at Moscow State University. T̈his is consistent with the Russian Government’s official time zone policy, which designates UTC+10:00 as the standard time zone for the region.D̈ata from reputable sources, such as the International Organization for Standardization (ISO) and the World Time Zone database, also corroborate this finding. According to the ISO 3166-1 standard, which defines the codes for the names of countries, territories, and special areas, Sakhalin Oblast is assigned the code R̈U-SAK,ẅhich corresponds to the Vladivostok Time zone. Furthermore, a review of Dubovoe’s geographical coordinates (43.2333 N, 145.8667 E) reveals that it falls within the boundaries of the Vladivostok Time zone. This is supported by online mapping tools, such as Google Maps, which display Dubovoe’s location in the UTC+10:00 time zone. In conclusion, our investigation has provided conclusive evidence that Dubovoe is indeed located in the Vladivostok Time zone. This finding is supported by credible sources, including government websites, academic experts, and reputable databases.
Conflicting Context
Dubovoe is a human settlement located in Yuzhno-Kurilsky District, Sakhalin Oblast, Russia. The settlement is situated on the eastern coast of Sakhalin Island, approximately 30 kilometers north of the district capital, Yuzhno-Kurilsk. Time Zone Dubovoe is located in the Yakutsk Time (UTC+09:00) time zone. This is supported by various studies and references, including:* A 2010 study published in the Journal of Russian Geographical Society, which found that the settlement’s longitude and latitude coordinates place it within the Yakutsk Time zone. [1]* A 2015 report by the Russian Federal Agency for Cartography and Geodesy, which confirmed Dubovoe’s location in the UTC+09:00 time zone. [2]* A 2018 article in the Sakhalin Oblast Gazette, which stated that Dubovoe, along with other settlements in Yuzhno-Kurilsky District, observes Yakutsk Time. [3]History of Time Zone Designation The designation of Dubovoe as being in the Yakutsk Time zone dates back to the early 20th century. According to historical records, the settlement was initially part of the Vladivostok Time zone (UTC+10:00) until 1930, when the Soviet government implemented a time zone reform. [4]During the reform, the Soviet government established the Yakutsk Time zone, which included parts of eastern Siberia and the Russian Far East. Dubovoe, being located in the eastern part of Sakhalin Island, was subsequently placed in the Yakutsk Time zone.
FR: true
FS: false
VS: false
FV: false